Null hypothesis: products equally

Alternative hypothesis: #2 sells cheaper than #1

H0: µ1 = µ2, Ha: µ1 > µ2

N=20

Please assist in finding the following:

random predicted outcome.

second outcome of 16/20 statistic

difference

sample difference

standard error

df

standardized statistics

p-value

standard deviation

standard error

confidence interval within 95%

product name stat 1 stat 2 Price Comparison
a $                         14.39 $          10.39 $                   4.00
b $                         14.39 $          10.94 $                   3.45
c $                           9.99 $            9.99 $                       -  
d $                           7.19 $            3.99 $                   3.20
e $                         14.40 $          12.24 $                   2.16
f $                           8.99 $            8.59 $                   0.40
g $                           3.95 $            3.95 $                       -  
h $                           9.89 $            8.50 $                   1.39
i $                         14.40 $          12.14 $                   2.26
j $                         17.96 $          10.79 $                   7.17
k $                         16.19 $          13.99 $                   2.20
l $                         15.30 $          10.09 $                   5.21
m $                           6.97 $            6.97 $                       -  
n $                         13.49 $            8.99 $                   4.50
o $                         15.29 $            9.99 $                   5.30
p $                         14.80 $          12.33 $                   2.47
o $                         13.01 $          13.01 $                       -  
p $                         15.26 $          14.26 $                   1.00
q $                         11.70 $          10.66 $                   1.04
r $                         13.30 $          12.07 $                   1.23
Answer #1
Sample #1 Sample #2 difference , Di =sample1-sample2 (Di - Dbar)²
14.39 10.39 4.00 2.73
14.39 10.94 3.45 1.21
9.99 9.99 0.00 5.52
7.19 3.99 3.20 0.72
14.4 12.24 2.16 0.04
8.99 8.59 0.40 3.80
3.95 3.95 0.00 5.52
9.89 8.5 1.39 0.92
14.4 12.14 2.26 0.01
17.96 10.79 7.17 23.24
16.19 13.99 2.20 0.02
15.3 10.09 5.21 8.19
6.97 6.97 0.00 5.52
13.49 8.99 4.50 4.63
15.29 9.99 5.30 8.71
14.8 12.33 2.47 0.01
13.01 13.01 0.00 5.52
15.26 14.26 1.00 1.82
11.7 10.66 1.04 1.71
13.3 12.07 1.23 1.25
Sample #1 Sample #2 difference , Di =sample1-sample2 (Di - Dbar)²
14.39 10.39 4.00 2.73
14.39 10.94 3.45 1.21
9.99 9.99 0.00 5.52
7.19 3.99 3.20 0.72
14.4 12.24 2.16 0.04
8.99 8.59 0.40 3.80
3.95 3.95 0.00 5.52
9.89 8.5 1.39 0.92
14.4 12.14 2.26 0.01
17.96 10.79 7.17 23.24
16.19 13.99 2.20 0.02
15.3 10.09 5.21 8.19
6.97 6.97 0.00 5.52
13.49 8.99 4.50 4.63
15.29 9.99 5.30 8.71
14.8 12.33 2.47 0.01
13.01 13.01 0.00 5.52
15.26 14.26 1.00 1.82
11.7 10.66 1.04 1.71
13.3 12.07 1.23 1.25
sample 1 sample 2 Di (Di - Dbar)²
sum = 250.86 203.88 46.980 81.080

Ho :   µd=   0                  
Ha :   µd >   0                  
                          
Level of Significance ,    α =    0.05       claim:µd=0          
                          
sample size ,    n =    20                  
                          
mean of sample 1,    x̅1=   12.543                  
                          
mean of sample 2,    x̅2=   10.194                  
                          
sample  difference ,    D̅ =ΣDi / n =   2.349                  
                          
std dev of difference , Sd =    √ [ (Di-Dbar)²/(n-1) =    2.0658                  
                          
std error , SE = Sd / √n =    2.0658   / √   20   =   0.4619      
                          
t-statistic = (D̅ - µd)/SE = (   2.349   -   0   ) /    0.4619   =   5.085
                          
Degree of freedom, DF=   n - 1 =    19                  
  
p-value =        0.000033   [excel function: =t.dist.rt(t-stat,df) ]               

sample size ,    n =    20          
Degree of freedom, DF=   n - 1 =    19   and α =    0.05  
t-critical value =    t α/2,df =    2.0930   [excel function: =t.inv.2t(α/2,df) ]      
                  
std dev of difference , Sd =    √ [ (Di-Dbar)²/(n-1) =    2.0658          
                  
std error , SE = Sd / √n =    2.0658   / √   20   =   0.4619
margin of error, E = t*SE =    2.0930   *   0.4619   =   0.9668
                  
mean of difference ,    D̅ =   2.349          
confidence interval is                   
Interval Lower Limit= D̅ - E =   2.349   -   0.9668   =   1.382
Interval Upper Limit= D̅ + E =   2.349   +   0.9668   =   3.316
                  
so, confidence interval is (   1.3822   < µd <   3.3158   )